3. Andrea wants to deposit money into a bank account that earns 4% simple annual interest. Use the formula to find the amount of money that she should deposit so that she earns $2800 after 3.5 years. Show and explain your work.
Answer:

Answer by stanbon(75887)

Andrea wants to deposit money into a bank account that earns 4% simple annual interest. Use the formula to find the amount of money that she should deposit so that she earns $2800 after 3.5 years.
-------
Formula: A(t) = P(1 + rt)
2800 = P(1 + 0.04*3.5)
2800 = P(1.14)
P = 2800/1.14 = $2456.14
---------------
Cheers,
Stan H.
---------------